OpenStudy (anonymous):

solve for x: 4x+6=x^2+5x please help !

OpenStudy (anonymous):

I will help you out.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i think x=2 but I'm not sure

OpenStudy (anonymous):

Ok, in order to solve this, it needs to be in standard form of a quadratic, this is something in the form \[ax^2+bx+c=0\] The first thing that you have to do with your equation is rearrange terms until it looks like that form. To do this, we subtract 4x and 6 from the left side. the result is: \[0=x^2-x-6\] Now we want to see if we can factor it, which is possible. You know it is going to be in the form of (x+ )(x- ). Factors of -6 that add up to -1??? , 3 and -2 work, so lets plug those in. The factored form is then (x+3)(x-2) The equation is now \[0=(x+3)(x-2)\] Which is solved when EITHER of the two quantities on the right hand side are zero (anything multiplied by zero is zero! This makes the equation true, and is a solution) This occurs when either (x+3) = 0 or when (x-2) = 0, so x must either be -3 or 2.

OpenStudy (anonymous):

Remember, you can always plug your solution back into the original equation to check it.

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

okay that makes sense i knew 2 was part of the answer but didn't know if that was the only answer..thanks so much for your help ! (:

OpenStudy (anonymous):

you are welcome, glad it helped